### v1

# Moment generating functions and moderate deviation principles for lacunary trigonometric sums

Abstract.
In a recent paper, Aistleitner, Gantert, Kabluchko, Prochno and Ramanan studied large deviation principles (LDPs) for lacunary trigonometric sums $\sum_{n=1}^{N}\cos(2\pi n_{k}x)$, where the sequence $(n_{k})_{k\geq 1}$ satisfies the Hadamard gap condition $n_{k+1}/n_{k}\geq q>1$ for $k\geq 1$. A crucial ingredient in their work were asymptotic estimates for the moment generating function (MGF) of such sums, which turned out to depend on the fine arithmetic structure of the sequence $(n_{k})_{k\geq 1}$ in an intricate way. In the present paper we carry out a detailed study of the MGF for lacunary trigonometric sums (without any structural assumptions on the underlying sequence, other than lacunarity), and we determine the sharp threshold where arithmetic effects start to play a role. As an application, we prove moderate deviation principles for lacunary trigonometric sums, and show that the tail probabilities are in accordance with Gaussian behavior throughout the whole range between the central limit theorem and the LDP regime.
Key words and phrases:
Lacunary trigonometric sums, Hadamard gap condition, moment generating function, moderate deviation principle
2020 Mathematics Subject Classification:
Primary 42A55; Secondary 11K06, 42A70, 60F05, 60F10
1. Introduction and main result
Lacunary trigonometric sums have a long and fascinating history. They are well-known for exhibiting properties which are fundamentally different from those of general trigonometric sums, and rather coincide in many cases with the behavior which is typically also observed for sums of independent random variables. Early examples of this phenomenon include Weierstraß’ construction of a continuous, nowhere differentiable function, and Fatou’s work on the almost everywhere convergence of lacunary trigonometric series, a topic which was brought to a culmination point by Kolmogorov [19]. Among the most remarkable early papers on lacunary sums, from a probabilistic perspective, are Salem and Zygmund’s [23] central limit theorem for lacunary trigonometric sums, Erdős and Gál’s [15] law of the iterated logarithm, and Kac’ work [18] on interference phenomena.
The most classical lacunary growth condition, called the Hadamard gap condition, requests that there exists a constant $q>1$ such that
(1) $\frac{n_{k+1}}{n_{k}}\geq q,\qquad k\geq 1,$
and for such a sequence of frequencies one studies the sums
$\sum_{k=1}^{N}\cos(2\pi n_{k}x)\qquad\text{and}\qquad\sum_{k=1}^{N}\sin(2\pi n_{k}x)$
(which behave very similarly in most regards, so that often the results and proofs are only stated for the cosine sums). Among possible generalizations, all of which have been studied extensively, are: the introduction of a coefficient sequence $(c_{k})_{k\geq 1}$, so that the sum becomes $\sum_{k=1}^{N}c_{k}\cos(2\pi n_{k}x)$, see for example Weiss [25] and Frühwirth and Hauke [16]; replacing the Hadamard gap condition by a weaker growth condition, see for example Erdős [14], Takahashi [24], Berkes [7], and Bobkov and Götze [10]; replacing $\cos(2\pi x)$ by a more general 1-periodic function $f(x)$, see for example Kac [18], Gapoškin [17], and Aistleitner and Berkes [1]; replacing the trigonometric system by another orthogonal system such as the Walsh system, see for example Levizov [20] and Berkes and Philipp [9]; taking a perspective from ergodic theory, see for example Conze and Le Borgne [11] and Conze, Le Borgne and Roger [12]; and interpreting lacunary sums in the wider framework of the “subsequence principle” of probability theory and functional analysis, see Aldous [5, 6] and Berkes and Péter [8]. For more general background on lacunary sums and related topics, we refer to the recent survey article [2].
It is notable that the central limit theorem and the law of the iterated logarithm for lacunary trigonometric sums hold in a universal form: if $(n_{k})_{k\geq 1}$ satisfies (1), then for all $t\in\mathbb{R}$
$\mathbb{P}\left(x\in(0,1):~{}\frac{1}{\sqrt{N}}\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\leq t\right)\to\Phi(t)\qquad\text{as $N\to\infty$,}$
where $\mathbb{P}$ is the Lebesgue measure and $\Phi$ is the distribution function of the standard normal distribution. Further,
[…]
the normalization factor $\sqrt{2}$, which appears in both formulas, comes from the standard deviation of the identically distributed (but not independent) random variables $\cos(2\pi n_{k}x),~{}k\geq 1,$ which is
$\left(\int_{0}^{1}\cos(2\pi n_{k}x)^{2}~{}\mathrm{d}x\right)^{1/2}=\frac{1}{\sqrt{2}}.$
We emphasize again that neither formula depends in any way on specific (for example, number-theoretic) properties of the sequence $(n_{k})_{k\geq 1}$, other than the assumption that it satisfies the growth condition (1). This “universal” behavior at the scales of central limit theorem and law of the iterated logarithm changes dramatically when $\cos(2\pi x)$ is replaced by a more general 1-periodic function $f(x)$, and many papers have been devoted to the intriguing effect that arithmetic properties of $(n_{k})_{k\geq 1}$ have on the probabilistic behavior of $\sum f(n_{k}x)$ for lacunary $(n_{k})_{k\geq 1}$; however, this is not the direction which we will pursue here, since throughout this paper we will stick to the case of pure trigonometric sums.
A very interesting effect was observed in a recent paper of Aistleitner, Gantert, Kabluchko, Prochno and Ramanan [4]. They studied so-called large deviation principles (LDPs) for lacunary trigonometric sums, which are essentially asymptotic estimates for the “large deviation” probabilities
$\mathbb{P}\left(x\in(0,1):~{}\frac{1}{N}\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)>t\right)$
(note that the scaling factor here is $1/N$, in contrast to the $1/\sqrt{N}$ of the central limit theorem). A comparison with the tail probabilities of the normal distribution suggests that these probabilities should be of order roughly $e^{-Nt^{2}/2}$, so it makes sense to consider the “rate function”
$I(t):=-\lim_{N\to\infty}\frac{1}{N}\log\left(\mathbb{P}\left(x\in(0,1):~{}\frac{1}{N}\sum_{k=1}^{N}\cos(2\pi n_{k}x)>t\right)\right),$
provided such a limit actually exists. By the Gärtner–Ellis theorem, the calculation of such rate functions is essentially reduced to the calculation of the moment generating function, or more precisely, the calculation of the limit as $N\to\infty$ of
$\Lambda_{N}(\lambda):=\frac{1}{N}\log\left(\mathbb{E}\left(\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\right)\right),$
provided this limit exists (throughout this paper, expected values are always understood to be taken with respect to the Lebesgue measure on $(0,1)$). Now, if the system $(\cos(2\pi n_{k}x))_{k\geq 1}$ were to exhibit truly independent behavior, then one could calculate
$\mathbb{E}\left(\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\right)=$ $\prod_{k=1}^{N}\mathbb{E}\left(\exp\left(\lambda\sqrt{2}\cos(2\pi n_{k}x)\right)\right)$
$=$ $\left(\mathbb{E}\left(\exp\left(\lambda\sqrt{2}\cos(2\pi x)\right)\right)\right)^{N},$
and accordingly one could expect the limit $\Lambda(\lambda)$ of $\Lambda_{N}(\lambda)$ to be
$\log\left(\int_{0}^{1}e^{\lambda\sqrt{2}\cos(2\pi x)}~{}\mathrm{d}x\right)=$ $\log\left(I_{0}\left(\sqrt{2}\lambda\right)\right)$
$=$ $\log\left(\sum_{m=0}^{\infty}\frac{\lambda^{2m}}{2^{m}(m!)^{2}}\right)$
(2) $=$ $\frac{\lambda^{2}}{2}-\frac{\lambda^{4}}{16}+\frac{\lambda^{6}}{72}-+\dots,$
where $I_{0}$ denotes a modified Bessel function of the first kind. Indeed, in [4] it is proved that under the stronger gap condition $\frac{n_{k+1}}{n_{k}}\to\infty$ as $k\to\infty$, the limiting function $\Lambda(\lambda)$ exists and indeed equals $I_{0}\left(\sqrt{2}\lambda\right)$, in accordance with true independent behavior. However, for other sequences $(n_{k})_{k\geq 1}$ (of actual exponential growth, and not super-exponential growth), the MGF of $\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)$ could only be dealt with in [4] with the necessary precision under strong structural assumptions on $(n_{k})_{k\geq 1}$, such as assuming the very special (“stationary”) situation $n_{k}=a^{k}$ for $k\geq 1$, for some integer $a\geq 2$.
In this paper, we study $\Lambda_{N}(\lambda)$ as a function of $\lambda\in[-1,1]$ for general lacunary sequences that satisfy (1). We prove (Theorem 1) that $\Lambda_{N}(\lambda)=\frac{\lambda^{2}}{2}+O(\lambda^{3})$ as $N\to\infty$ throughout the range $\lambda\in[-1,1]$ for all lacunary sequences $(n_{k})_{k\geq 1}$. Thus the quadratic term in the expansion of $\Lambda_{N}(\lambda)$ is always (asymptotically) in accordance with the random model, and also in accordance with Gaussian behavior. However, the cubic term does not need to be in agreement with the random model, and (as the examples after the statement of the theorem will show) our error term $O(\lambda^{3})$ is optimal as far as general lacunary sequences are considered.
Theorem 1.
Let $(n_{k})_{k\geq 1}$ be a sequence of positive integers satisfying the Hadamard gap condition $\frac{n_{k+1}}{n_{k}}\geq q$, $k\geq 1$, for some constant $q>1$. Then for all $\lambda\in[-1,1]$ we have
$\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x=\exp\left(\frac{\lambda^{2}N}{2}+O\left(\lambda^{3}N\right)\right)$
as $N\to\infty$, where the implied constant depends only on $q$. In particular, we have
$\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x=\exp\left(\frac{\lambda^{2}N}{2}(1+o(1))\right)$
as $N\to\infty$, provided that $\lambda=o(1)$.
We give a few examples, which show that the term $O(\lambda^{3})$ in the statement of the theorem is optimal.
  • The sequence $n_{k}=2^{k},~{}k\geq 1$. This is the sequence which was already studied by Kac [18]. Note that this is a structurally very special case, since it has a very clear interpretation in terms of ergodic theory, where it can be seen as the action of the doubling map $T:~{}x\mapsto 2x$ mod $1$. In this case a limit function $\Lambda(\lambda)=\lim_{N\to\infty}\Lambda_{N}(\lambda)$ indeed exists, and is calculated (using combinatorial tools) in the appendix of [4], where it is shown to be
    $\Lambda(\lambda)=\frac{\lambda^{2}}{2}+\frac{\lambda^{3}}{2\sqrt{2}}+\frac{3\lambda^{4}}{16}+\dots.$
    Note the presence of a cubic term, which is absent in the series (2) for the random model.
  • We remark that the calculation of $\Lambda(\lambda)$ for $n_{k}=2^{k},~{}k\geq 1$ is somewhat involved, since the system $\cos(2\pi 2^{k}x))_{k\geq 1}$ has a rather complicated dependence structure. A simpler model is $n_{1}=1$ and
    $n_{k+1}=2n_{k}\text{ for $k$ odd,}\qquad n_{k+1}=k!n_{k}\text{ for $k$ even}.$
    Here, the factor $k!$ is quite arbitrarily chosen, since it is only used to break the overall dependence structure into pairs of consecutive indices. Using the fact that $n_{2k+1}/n_{2k}$ increases very quickly, one can show for this sequence that
    $\mathbb{E}\left(\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\right)\approx\left(\mathbb{E}\left(\exp\left(\lambda\sqrt{2}\left(\cos(2\pi x)+\cos(4\pi x)\right)\right)\right)\right)^{N/2}$
    as $N\to\infty$, and thus
    $\Lambda(\lambda)=\frac{1}{2}\log\left(\int_{0}^{1}\exp\left(\lambda\sqrt{2}\left(\cos(2\pi x)+\cos(4\pi x)\right)\right)\mathrm{d}x\right)=\frac{\lambda^{2}}{2}+\frac{\lambda^{3}}{4\sqrt{2}}-\frac{\lambda^{4}}{16}+\dots$
    Note again the presence of the cubic term.
  • Another simple model is the sequence with $n_{1}=1$, and
    $n_{k+1}=2n_{k}\quad\text{and}\quad n_{k+2}=3n_{k}\qquad\text{for $k\equiv 1$ mod 3},$
    and $n_{k}=k!n_{k-1}$ when $k\equiv 0$ mod 3. The dependence structure of this sequence decomposes into triples of consecutive indices, and one has
    $\mathbb{E}\left(\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\right)$
    $\approx$ $\left(\mathbb{E}\left(\exp\left(\lambda\sqrt{2}\left(\cos(2\pi x)+\cos(4\pi x)+\cos(6\pi x)\right)\right)\right)\right)^{N/3}$
    as $N\to\infty$, and thus
    $\Lambda(\lambda)=$ $\frac{1}{3}\log\left(\int_{0}^{1}\exp\left(\lambda\sqrt{2}\left(\cos(2\pi x)+\cos(4\pi x)+\cos(6\pi x)\right)\right)\mathrm{d}x\right)$
    $=$ $\frac{\lambda^{2}}{2}+\frac{\lambda^{3}}{2\sqrt{2}}+\frac{7\lambda^{4}}{144}+\dots$
    Note again the presence of the cubic term, with yet another coefficient. We further remark that, from a perspective of Diophantine equations (which is from a technical point of view the key to understanding lacunary sums), this sequence is rather different from the two sequences mentioned before. In the above cases, the sequences were constructed in such a way that there are many solutions $k$ of the equation $n_{k+1}=2n_{k}$, and the reason why these solutions contribute to the cubic term in the expansion of $\lambda$ is that this 2-term Diophantine equation has coefficients $1$ and $2$, such that the sum of coefficients is $1+2=3$, with 3 meaning “cubic” (the way how the coefficients of these Diophantine equations arise from powers of $\lambda\cos(2\pi n_{k}x)$, and the total power of $\lambda$ is thus related to the sum of the coefficients of the Diophantine equations, will be clearly visible in the proof of Theorem 1). Now in this new example, the sequence is constructed in such a way that there are many solutions $k$ of $n_{k+1}=n_{k}+n_{k-1}$, which is a 3-term Diophantine equation with coefficients $1,1,1$, leading to a non-vanishing cubic term in the expansion of $\Lambda(\lambda)$ since 1+1+1=3. The sequence in this last example is modeled on the Fibonacci sequence $(F_{k})_{k\geq 1}$, which of course also satisfies $F_{k+2}=F_{k+1}+F_{k}$ (for all $k$, not only for those congruent 0 mod 3). One could probably also study the Fibonacci sequence itself, rather than the sequence $(n_{k})_{k\geq 1}$ as constructed above, and prove that $\Lambda_{N}(\lambda)$ converges to some limiting function $\Lambda(\lambda)$ with non-vanishing cubic term in its series expansion, but the dependence structure of $(\cos(2\pi F_{k}x))_{k\geq 1}$ is much more delicate than that of our construction, and it is a priori not even clear that $\Lambda_{N}(\lambda)$ converges for $(F_{k})_{k\geq 1}$. Note that in contrast to $(\cos(2\pi 2^{k}x))_{k\geq 1}$, the system $(\cos(2\pi F_{k}x))_{k\geq 1}$ does not seem to have an immediate interpretation from the perspective of ergodic theory.
  • From the discussion above, it is not difficult to construct a lacunary sequence $(n_{k})_{k\geq 1}$ for which $\Lambda_{N}(\lambda)$ fluctuates and does not converge as $N\to\infty$ (beyond the quadratic term in the series expansion). Such a sequence could for example be constructed by “mixing” finite segments (of increasing length) of the sequences from the previous examples. Together with the examples above, this shows that indeed $\Lambda_{N}(\lambda)=\lambda^{2}/2+O(\lambda^{3})$ is the best one could hope for, for general lacunary sequences $(n_{k})_{k\geq 1}$.
We note in passing that our method of proof entails a potential dependence on the growth factor $q$ for the implied constant in the $O(\lambda^{3})$ term, in the sense that this implied constant possibly needs to be increased as $q\to 1$. It seems plausible that this $q$-dependent implied constant could actually be replaced by an absolute constant, the optimal value for such a constant possibly being $1/\sqrt{2}$ (arising from the case of the Fibonacci sequence). However, we have not pursued this topic any further.
Our ability to provide optimal estimates for the (asymptotic) MGF of $\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)$ allows us to establish moderate deviation principles (MDPs) (see Theorem 2) for the latter via the Gärtner-Ellis theorem (see, e.g., [13, Theorem 2.3.6]). These principles, originating from [22], have become an important concept in the study of the distribution of sequences of random variables. Let $\lambda=\lambda_{N}$ be a null sequence with $\lambda\sqrt{N}\to\infty$ and let $(X_{N})_{N\in\mathbb{N}}$ be a sequence of random variables defined on $(0,1)$. We say that $\left(X_{N}\right)_{N\in\mathbb{N}}$ satisfies an MDP at speed $1/\lambda^{2}$ if there exists $\sigma^{2}>0$ such that for all Borel-sets $B\subseteq\mathbb{R}$ we have
(3) $-\inf_{x\in B^{\circ}}\frac{x^{2}}{2\sigma^{2}}\leq\liminf_{N\rightarrow\infty}\lambda^{2}\log\mathbb{P}\left(X_{N}\in B\right)\leq\limsup_{N\rightarrow\infty}\lambda^{2}\log\mathbb{P}\left(X_{N}\in B\right)\leq-\inf_{x\in\overline{B}}\frac{x^{2}}{2\sigma^{2}},$
where $B^{\circ}$ is the interior and $\overline{B}$ is the closure of $B$. If one considers a sum of independent random variables $Y_{1}+\ldots+Y_{N}$, all having mean 00 and variance $\sigma^{2}$, it is well-known that $X_{N}:=\frac{\lambda}{\sqrt{N}}\sum_{k=1}^{N}Y_{k}$ satisfies an MDP at speed $1/\lambda^{2}$ with variance $\sigma^{2}$ (see, e.g., [13, Theorem 3.7.1]). Heuristically speaking, for such a sequence, the central limit theorem would suggest
$\mathbb{P}\left(\frac{\lambda}{\sqrt{N}}\sum_{k=1}^{N}Y_{k}\geq x\right)\approx e^{-\frac{x^{2}}{2(\lambda\sigma)^{2}}},\quad\text{as}\quad N\longrightarrow\infty.$
An MDP tells us that, on a logarithmic scale, the above holds even for arbitrary Borel sets instead of intervals of the form $[x,\infty)$ for some $x>0$. Our Theorem 1 allows us to obtain very general moderate deviation principles for trigonometric lacunary sums, without the need for any additional assumptions other than lacunarity.
Theorem 2.
Let $(n_{k})_{k\in\mathbb{N}}$ be a lacunary sequence satisfying (1). Then, for any positive null sequence $\lambda=\lambda_{N}$ with $\sqrt{N}\lambda\to\infty$,
$\frac{\lambda}{\sqrt{N}}\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)$
satisfies an MDP at speed $1/\lambda^{2}$ in the sense of (3) with $\sigma^{2}=1$.
One should compare Theorem 2, which gives a clean universal result, to the results in [21], where MDPs for lacunary sums $\sum f(n_{k}x)$ generated by trigonometric polynomials $f$ were studied. There, the authors had to impose strong arithmetic assumptions on the lacunary sequence $(n_{k})_{k\geq 1}$, as well as an additional growth condition upon $\lambda$, in order to obtain their moderate deviation principles (see [21, Theorem 1]). It is not surprising that stronger assumptions are necessary to obtain MDPs in the general case $\sum f(n_{k}x)$, since CLT and LIL also hold in the case of $\sum f(n_{k}x)$ only under additional arithmetic assumptions. However, [21] provides only sufficient criteria for the validity of an MDP, while sharp (necessary and sufficient) arithmetic conditions are known in the CLT and LIL cases, see [1, 3].
2. Proofs
Proof of Theorem 1.
Assume that $N$ is “large”, and let $\lambda\in\mathbb{R}$ be given. Let $s$ (“short”) and $L$ (“long”) be two positive integers, to be specified later. We decompose the set of indices $\{1,2,\dots,N\}$ into a disjoint union of blocks
(4) $\{1,\dots,N\}=\Delta_{1}\cup\Delta_{1}^{\prime}\cup\Delta_{2}\cup\Delta_{2}^{\prime}\cup\cdots\cup\Delta_{M}\cup\Delta_{M}^{\prime},$
such that $|\Delta_{i}|=L$ and $|\Delta_{i}^{\prime}|=s$ for $1\leq i\leq M$, where $M=N/(L+s)$. (To simplify notations we assume that indeed $N$ is divisible by $s+L$ without remainder; otherwise one has to shorten one of the last two blocks and possibly skip $\Delta_{M}^{\prime}$, which does not affect the overall calculation.) The decomposition in (4) has to be understood in the sense that
$\Delta_{1}<\Delta_{1}^{\prime}<\Delta_{2}<\Delta_{2}^{\prime}<\dots$
holds element-wise. The sum over all the long blocks $\Delta_{i}$ will give the main contribution, and the role of the short blocks $\Delta_{i}^{\prime}$ is to ensure that the frequencies $n_{k}$ and $n_{\ell}$ of two trigonometric functions $\cos(2\pi n_{k}x)$ and $\cos(2\pi n_{\ell}x)$ with indices $k$ and $\ell$ coming from two different long blocks are of very different size, thus creating a high degree of “independence” between trigonometric functions (interpreted as random variables) whenever the respective indices are in different long blocks.
Recall that $q>1$ is the growth factor of the lacunary sequence under consideration. We define $s$ as the smallest positive integer for which $\min(q-1,1/2)-4\frac{q^{-1}}{q^{s}-1}>0$; note that $s$ depends on $q$, but clearly not on $N$ or $\lambda$. Furthermore, we set $L=\left\lceil\frac{1}{5|\lambda|}\right\rceil$. Note that the conclusion of the theorem is trivial unless $|\lambda|\to 0$ as $N\to\infty$. Thus we can assume, throughout the rest of the paper, that $N$ and $\lambda$ are such that
(5) $s<L\qquad\text{and}\qquad L\leq\frac{1}{4|\lambda|}.$
We have
(6) $\begin{split}&\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x\\ =&\int_{0}^{1}\prod_{i=1}^{M}\exp\left(\lambda\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)\cdot\prod_{i=1}^{M}\exp\left(\lambda\sum_{k\in\Delta_{i}^{\prime}}\sqrt{{2}}\cos(2\pi n_{k}x)\right)\mathrm{d}x,\end{split}$
and the aim now is to estimate this integral of a product in terms of a product of integrals (note that under actual stochastic independence, the integral of a product would automatically decompose into a product of integrals, so what we are trying to achieve here is to mimic such independent behavior).
We write $k_{i,1},\dots,k_{i,s}$ for the indices in $\Delta_{i}^{\prime}$, sorted in increasing order. Using $e^{x}=e^{x^{2}/2}e^{x-x^{2}/2}$ and writing $g(x)=e^{x-x^{2}/2}$, we get
$\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x$
$=$ $\,\int_{0}^{1}\prod_{i=1}^{M}\exp\left(\lambda\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)\cdot\prod_{j=1}^{s}\prod_{i=1}^{M}\exp\left(\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)\mathrm{d}x$
(7) $\leq$ $\,e^{\lambda^{2}sM}\int_{0}^{1}\prod_{i=1}^{M}\exp\left(\lambda\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)\cdot\prod_{j=1}^{s}\prod_{i=1}^{M}g\left(\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)\mathrm{d}x.$
We apply Hölder’s inequality for a product of $s+1$ functions, with the first exponent being $p=\frac{L+s}{L}<2$, and all the remaining exponents being $p^{\prime}=L+s$. Thus (7) is bounded above by
(8) $e^{\lambda^{2}sM}I_{M}(\lambda)^{1/p}\cdot\prod_{j=1}^{s}\left(\int_{0}^{1}\prod_{i=1}^{M}g\left(\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)^{p^{\prime}}\mathrm{d}x\right)^{\frac{1}{p^{\prime}}},$
where
(9) $I_{M}(\lambda):=\int_{0}^{1}\prod_{i=1}^{M}\exp\left(p\lambda\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x.$
Next, we will provide an upper bound for $I_{M}(\lambda)$. One can easily show that $0<\exp(x)\leq 1+x+\frac{x^{2}}{2}+\frac{x^{3}}{6}+x^{4}$ for $|x|\leq 1$. Since by (5) we have
(10) $\left|p\lambda\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right|\leq\sqrt{2}p|\lambda|L=2\sqrt{2}|\lambda|\frac{1}{4|\lambda|}\leq 1,$
we obtain
$\exp\left(p\lambda\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)$
$\leq$  $\,1+\lambda p\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)+\frac{(\lambda p)^{2}}{2}\left(\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)^{2}$
$\phantom{\,1}+\frac{(\lambda p)^{3}}{6}\left(\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)^{3}+(\lambda p)^{4}\left(\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)^{4}.$
We will expand these sums, keeping special track of the diagonal terms in the squared expression, for which we use $\cos(x)^{2}=1/2+\cos(2x)/2$. Hence we can rewrite the expression in the last two displayed lines as
$S_{i}(x):=$  $\,1+\frac{\lambda^{2}p^{2}L}{2}+\lambda p\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)+\frac{\lambda^{2}p^{2}}{2}\sum_{k\in\Delta_{i}}\cos(2\pi 2n_{k}x)$
$\qquad+\lambda^{2}p^{2}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\Delta_{i},\\ k_{1}\neq k_{2}\end{subarray}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)$
$\qquad+\frac{\lambda^{3}p^{3}2^{3/2}}{6}\sum_{k_{1},k_{2},k_{3}\in\Delta_{i}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)\cos(2\pi n_{k_{3}}x)$
$\qquad+4\lambda^{4}p^{4}\sum_{k_{1},k_{2},k_{3},k_{4}\in\Delta_{i}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)\cos(2\pi n_{k_{3}}x)\cos(2\pi n_{k_{4}}x).$
Recall that $\cos(x)\cos(y)=\frac{\cos(x+y)}{2}+\frac{\cos(x-y)}{2}$. Accordingly, all the products of cosine-functions which appear in the definition of $S_{i}(x)$ can be rewritten into cosines of sums/differences of frequencies such as $n_{k_{1}}\pm n_{k_{2}}$, $n_{k_{1}}\pm n_{k_{2}}\pm n_{k_{3}}$ or $n_{k_{1}}\pm n_{k_{2}}\pm n_{k_{3}}\pm n_{k_{4}}$. We will denote by $i_{-}$ and $i_{+}$ the smallest and the largest elements of $\Delta_{i}$, respectively. Then the sum $\sum_{k\in\Delta_{i}}\cos(2\pi n_{k}x)$ in the definition of $S_{i}(x)$ is a trigonometric polynomial with frequencies between $n_{i_{-}}$ and $n_{i_{+}}$. Likewise, the sum $\sum_{k\in\Delta_{i}}\cos(2\pi 2n_{k}x)$ is a trigonometric polynomial with frequencies between $2n_{i_{-}}$ and $2n_{i_{+}}$. For the next term, we use the fact that our sequence $(n_{k})_{k\geq 1}$ is of lacunary growth, to note that
$T_{i}^{(2)}(x)$ $:=\lambda^{2}p^{2}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\Delta_{i},\\ k_{1}\neq k_{2}\end{subarray}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)$
$\phantom{:}=\lambda^{2}p^{2}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\Delta_{i},\\ k_{1}\neq k_{2}\end{subarray}}\frac{\cos(2\pi(n_{k_{1}}+n_{k_{2}})x)}{2}+\frac{\cos(2\pi(n_{k_{1}}-n_{k_{2}})x)}{2}$
is a trigonometric polynomial with frequencies between $(q-1)n_{i_{-}}$ and $2n_{i_{+}}$. The following terms are a bit more difficult to analyze.
Consider $\sum_{k_{1},k_{2},k_{3}\in\Delta_{i}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)\cos(2\pi n_{k_{3}}x)$. The product of cosines can be re-written as $\cos(2\pi(\pm n_{k_{1}}\pm n_{k_{2}}\pm n_{k_{3}}))$. If all three “$\pm$” signs are the same, then such a cosine has a frequency between $3n_{i_{-}}$ and $3n_{i_{+}}$. For other combinations of signs, we are interested in how often the frequency |±$|\pm n_{k_{1}}\pm n_{k_{2}}\pm n_{k_{3}}|$ can be “small”. By symmetry, it suffices to count the number of such triples of indices in the case where the combinations of signs is “(+,,)(+,-,-)( + , - , - )”. Thus, let us assume that
(11) $\left|n_{k_{1}}-n_{k_{2}}-n_{k_{3}}\right|<\frac{n_{i_{-}}}{2}$
for some $(k_{1},k_{2},k_{3})\in\Delta_{i}^{3}$, and see how often this is possible. Assume w.l.o.g. that $k_{2}\geq k_{3}$. If $k_{2}\geq k_{1}$, then $n_{k_{1}}-n_{k_{2}}-n_{k_{3}}\leq-n_{k_{3}}\leq-n_{i_{-}}$, so in this case (11) fails. Thus $k_{1}>k_{2}$ (recall that we also assumed $k_{2}\geq k_{3}$). By the lacunary growth condition, there are at most $O(1)$ many choices of $k_{2}$ with $k_{2}<k_{1}$ such that $n_{k_{2}}\geq\frac{n_{k_{1}}}{4}$, which is easily seen to be a necessary condition for (11). For given $k_{1}$ and $k_{2}$, there are at most $O(1)$ many values of $k_{3}$ for which (11) holds, again as a consequence of the lacunary growth condition. Overall, that means that the number of triples $(k_{1},k_{2},k_{3})\in\Delta_{i}^{3}$ for which (11) can hold is of cardinality at most $O(L)$. In other words, we can write
$\frac{\lambda^{3}p^{3}2^{3/2}}{6}\sum_{k_{1},k_{2},k_{3}\in\Delta_{i}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)\cos(2\pi n_{k_{3}}x)=T_{i}^{(3)}(x)+O(\lambda^{3}L),$
where $T_{i}^{(3)}(x)$ is a suitable trigonometric polynomial whose frequencies are all between $\frac{n_{i_{-}}}{2}$ and $3n_{i_{+}}$.
Now we carry out a similar analysis for
$\sum_{k_{1},k_{2},k_{3},k_{4}\in\Delta_{i}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)\cos(2\pi n_{k_{3}}x)\cos(2\pi n_{k_{4}}x).$
This can be re-written into a sum of cosine-functions of the form $\cos(2\pi(\pm n_{k_{1}}\pm n_{k_{2}}\pm n_{k_{3}}\pm n_{k_{4}})$. If all four “$\pm$” signs are the same, then the frequency is between $4n_{i_{-}}$ and $4n_{i_{+}}$. The other two cases (up to symmetry) are those of signature (+,+,,)(+,+,-,-)( + , + , - , - ) and of signature (+,+,+,)(+,+,+,-)( + , + , + , - ).
Assume first that we study the case of signature (+,+,,)(+,+,-,-)( + , + , - , - ), and that for some $(k_{1},k_{2},k_{3},k_{4})\in\Delta_{i}^{4}$ we have
(12) $|n_{k_{1}}+n_{k_{2}}-n_{k_{3}}-n_{k_{4}}|<\frac{n_{i_{-}}}{2}.$
Assume w.l.o.g. that $k_{1}\geq k_{2}$, that $k_{3}\geq k_{4}$, and that $k_{1}\geq k_{3}$. It is easily seen that $n_{k_{3}}\geq\frac{n_{k_{1}}}{4}$ is a necessary condition for (12), since otherwise
$n_{k_{1}}+n_{k_{2}}-n_{k_{3}}-n_{k_{4}}\geq n_{k_{1}}-2n_{k_{3}}\geq\frac{n_{k_{1}}}{2}\geq\frac{n_{i_{-}}}{2}.$
Note that for given $k_{1}$, there are only $O(1)$ many choice of $k_{3}$ for which $k_{3}\leq k_{1}$ and $n_{k_{3}}\geq\frac{n_{k_{1}}}{4}$. Now, for any given $k_{1},k_{2},k_{3}$, there are at most $O(1)$ many choices of $k_{4}$ for which (12) holds. Thus, roughly speaking, in (12) the values of $k_{1},k_{2}$ are freely chosen (which we can do in $L^{2}$ many ways), which then restricts $k_{3},k_{4}$ to finitely many possibilities, and so the number of 4-tuples $(k_{1},k_{2},k_{3},k_{3})$ for which (12) holds is of order $O(L^{2})$.
Next we study the case of signature (+,+,+,)(+,+,+,-)( + , + , + , - ), and assume that for some $(k_{1},k_{2},k_{3},k_{4})\in\Delta_{i}^{4}$ we have
(13) $|n_{k_{1}}+n_{k_{2}}+n_{k_{3}}-n_{k_{4}}|<\frac{n_{i_{-}}}{2}.$
It is clear that this can only be true if $k_{4}>\max(k_{1},k_{2},k_{3})$. Assume w.l.o.g. that $k_{3}\geq\max(k_{1},k_{2})$. It is easily seen that $n_{k_{3}}\geq\frac{n_{k_{4}}}{6}$ is necessary for (13) to hold. Thus for given $k_{4}$, there are $O(1)$ many possible values for $k_{3}$. Furthermore, for given $k_{2}$, there are at most $O(1)$ many possible values for $k_{1}$. Thus the number of 4-tuples $(k_{1},k_{2},k_{3},k_{3})$ for which (13) holds is again of order $O(L^{2})$. Consequently, we can write
$4\lambda^{4}p^{4}\sum_{k_{1},k_{2},k_{3},k_{4}\in\Delta_{i}}\cos(2\pi n_{k_{1}}x)\cos(2\pi n_{k_{2}}x)\cos(2\pi n_{k_{3}}x)\cos(2\pi n_{k_{4}}x)$
$=$ $T_{i}^{(4)}(x)+\underbrace{O(\lambda^{4}L^{2})}_{=O(\lambda^{3}L)},$
where $T_{i}^{(4)}(x)$ is a trigonometric polynomial whose frequencies are all between $\frac{n_{i_{-}}}{2}$ and $4n_{i_{+}}$.
Combining all our findings, we have shown that there exists a trigonometric polynomial
$T_{i}(x):=\lambda p\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)+\frac{\lambda^{2}p^{2}}{2}\sum_{k\in\Delta_{i}}\cos(2\pi 2n_{k}x)+T_{i}^{(2)}(x)+T_{i}^{(3)}(x)+T_{i}^{(4)}(x),$
all of whose frequencies are between $\min(q-1,1/2)n_{i_{-}}$ and $4n_{i_{+}}$, such that
$S_{i}(x)=1+\frac{\lambda^{2}p^{2}L}{2}+T_{i}(x)+O(\lambda^{3}L).$
By construction, $i_{-}-(i-1)_{+}\geq s+1$ (since the long blocks are separated by short blocks of length $s$), which implies that
$n_{i_{-}}\geq q^{s+1}n_{(i-1)_{+}}.$
Similarly, we have $n_{i_{+}}\geq q^{L-1}n_{i_{-}}$, and combining this we obtain a chain of inequalities
(14) $n_{i_{-}}\geq q^{s+1}n_{(i-1)_{+}}\geq q^{2s+L+1}n_{(i-2)_{+}}\geq q^{3s+2L+1}n_{(i-3)_{+}}\geq\dots$
By the orthogonality properties of the trigonometric system, for any $m$ and any integers $\omega_{1}<\dots<\omega_{m}$, we have
(15) $\int_{0}^{1}\prod_{j=1}^{m}\cos(2\pi\omega_{j}x)\mathrm{d}x=0$
if there is there is no solution to $\omega_{m}\pm\omega_{m-1}\pm\cdots\pm\omega_{1}=0$. Thus, let us take a set of indices $i_{1}<\dots<i_{m}$ in $\{1,\dots,M\}$ and a set of frequencies $\{\omega_{j}\}_{j=1}^{m}$, such that $\omega_{j}$ is a frequency of one of the trigonometric functions that appear as a summand of $T_{i_{j}}$, for $1\leq j\leq m$. We will show for all possibilities of the $\pm$ signs, the expression $\omega:=\omega_{m}\pm\omega_{m-1}\pm\cdots\pm\omega_{1}$ is always positive, and thus the integral of the product vanishes as in (15). By
$\omega_{j}\in\left[\min\left(q-1,\frac{1}{2}\right)n_{{i_{j}}_{-}},4n_{{i_{j}}_{+}}\right]\quad\text{for all $j$},$
together with (14), we have
$\omega$ $\geq\omega_{m}-\omega_{m-1}-\cdots-\omega_{1}$
$\geq\min(q-1,1/2)n_{{i_{m}}_{-}}-4n_{{i_{m-1}}_{+}}-\cdots-4n_{{i_{1}}_{+}}$
$\geq\min(q-1,1/2)n_{{i_{m}}_{-}}-4q^{-s-1}n_{{i_{m}}_{-}}-4q^{-2s-L-1}n_{{i_{m}}_{-}}-\cdots-4q^{-(m-1)s-(m-2)L-1}n_{{i_{m}}_{-}}$
$=n_{{i_{m}}_{-}}\left(\min(q-1,1/2)-4q^{-s-1}\sum_{j=0}^{m-2}\frac{1}{(q^{s+L})^{j}}\right).$
Note that
$\min(q-1,1/2)-4q^{-s-1}\sum_{j=0}^{m-2}\frac{1}{(q^{s+L})^{j}}\geq$ $\min(q-1,1/2)-4q^{-s-1}\sum_{j=0}^{\infty}\frac{1}{(q^{s+L})^{j}}$
$=$ $\min(q-1,1/2)-4\frac{q^{L-1}}{q^{s+L}-1}.$
The function $\frac{x}{x-1}$ is decreasing on $(1,\infty)$, hence
$\min(q-1,1/2)-4\frac{q^{L-1}}{q^{s+L}-1}>\min(q-1,1/2)-4\frac{q^{-1}}{q^{s}-1}>0,$
where the last inequality follows from the choice of $s$.
This implies that for all $I\subseteq\{1,\ldots,M\}$,
$\int_{0}^{1}\prod_{i\in I}(1+T_{i}(x))\,\mathrm{d}x=1.$
Since $S_{i}(x)>0$ for all $x$, we accordingly obtain
$I_{M}(\lambda)=\int_{0}^{1}\prod_{i=1}^{M}S_{i}(x)\mathrm{d}x=$ $\int_{0}^{1}\prod_{i=1}^{M}\left(1+\frac{\lambda^{2}p^{2}L}{2}+T_{i}(x)+O(\lambda^{3}L)\right)\mathrm{d}x$
$=$ $\sum_{I\subseteq\{1,\ldots,M\}}\left(\int_{0}^{1}\prod_{i\in I}(1+T_{i}(x))\,\mathrm{d}x\right)\cdot\prod_{i\in\{1,\ldots,M\}\setminus I}\left(\frac{\lambda^{2}p^{2}L}{2}+O(\lambda^{3}L)\right)$
(16) $=$ $\prod_{i=1}^{M}\left(1+\frac{\lambda^{2}p^{2}L}{2}+O(\lambda^{3}L)\right)$
$\leq$ $\exp\left(\frac{\lambda^{2}p^{2}ML}{2}+O(\lambda^{3}ML)\right)$
(17) $\leq$ $\exp\left(\frac{\lambda^{2}p^{2}N}{2}+O(\lambda^{3}N)\right)$
since $M(L+s)=N$ and thus $ML\leq N$.
In view of (8), we are left to provide an upper bound for
(18) $\prod_{j=1}^{s}\left(\int_{0}^{1}\prod_{i=1}^{M}g\left(\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)^{p^{\prime}}\mathrm{d}x\right)^{\frac{1}{p^{\prime}}}.$
As a consequence of (5),
$|p^{\prime}\lambda\sqrt{2}\cos(2\pi x)|\leq\sqrt{2}|\lambda|(L+s)\leq 2|\lambda|\frac{1}{4|\lambda|}\leq 1/2,$
so we have $g\left(\lambda\sqrt{{2}}\cos(2\pi x)\right)^{p^{\prime}}\leq g(2p^{\prime}\lambda\sqrt{{2}}\cos(2\pi x))$, thus (18) is bounded above by
$\prod_{j=1}^{s}\int_{0}^{1}\prod_{i=1}^{M}g\left(2p^{\prime}\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\mathrm{d}x\right)^{\frac{1}{p^{\prime}}}.$
In the sequel, let $j$ be a fixed number from the set $\{1,\dots,s\}$. Using
$g(x)=e^{x-x^{2}/2}\leq\left(1+x-\frac{x^{3}}{3}\right)\quad\text{for}\quad|x|\leq 1,$
together with $|2p^{\prime}\lambda\sqrt{2}\cos(2\pi x)|\leq 1$ we have
$\prod_{i=1}^{M}g\left(2p^{\prime}\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)$
$\leq$ $\prod_{i=1}^{M}\left(1+2p^{\prime}\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)-\frac{\left(2p^{\prime}\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)^{3}}{3}\right)$
$=$ $\prod_{i=1}^{M}\left(1+\frac{4\sqrt{2}p^{\prime}\lambda-\sqrt{2}(2p^{\prime})^{3}\lambda^{3}}{2}\cos(2\pi n_{k_{i,j}}x)-\frac{\sqrt{2}(2p^{\prime})^{3}\lambda^{3}}{6}\cos(2\pi 3n_{k_{i,j}}x)\right).$
Here we used $\cos(x)^{3}=\frac{3}{4}\cos(x)+\frac{1}{4}\cos(3x)$ for the last line. Now we can use an orthogonality argument similar to the one starting in line (15), but way less delicate since the ratios $n_{k_{i+1,j}}/n_{k_{i,j}}$ are extremely large (the indices considered here are separated by long, not only by short blocks), to show that after multiplying out the product in the last displayed equation, everything non-constant vanishes after integration, and thus (for $N$ sufficiently large)
$\int_{0}^{1}\prod_{i=1}^{M}\left(1+\frac{2\sqrt{2}p^{\prime}\lambda-\sqrt{2}(p^{\prime})^{3}\lambda^{3}}{2}\cos(2\pi n_{k_{i,j}}x)-\frac{\sqrt{2}(p^{\prime})^{3}\lambda^{3}}{6}\cos(2\pi 3n_{k_{i,j}}x)\right)\mathrm{d}x=1,$
which proves
(19) $\int_{0}^{1}\prod_{i=1}^{M}g\left(p^{\prime}\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)\mathrm{d}x\leq 1.$
Combining this with (7),(8), and (17), and using $M=O(\lambda N)$ and $1/L=O(\lambda)$ we arrive at
$\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x\leq$ $\,e^{\lambda^{2}Ms}I_{M}^{1/p}$
$\leq$ $\,\exp\left(\frac{\lambda^{2}pN}{2}+\lambda^{2}Ms+O(\lambda^{3}N)\right)$
$=$ $\,\exp\left(\frac{\lambda^{2}N}{2}\left(1+\frac{s}{L}\right)+\lambda^{2}Ms+O(\lambda^{3}N)\right)$
(20) $=$ $\,\exp\left(\frac{\lambda^{2}N}{2}+O(\lambda^{3}N)\right).$
We are left to provide a matching lower bound which we will prove in a very similar way, using slightly different (Taylor) polynomial approximations for the exponential functions, and the reverse Hölder inequality. Similar to above, using $e^{x}\geq e^{x-x^{2}/2}=g(x)$, we now have
$\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x$
$\geq$ $\int_{0}^{1}\prod_{i=1}^{M}\exp\left(\lambda\sum_{k\in\Delta_{i}}\sqrt{2}\cos(2\pi n_{k}x)\right)\cdot\prod_{j=1}^{s}\prod_{i=1}^{M}g\left(\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)\mathrm{d}x$
(21) $\geq$ $I_{M}(\lambda)^{1/p}\cdot\prod_{j=1}^{s}\left(\int_{0}^{1}\prod_{i=1}^{M}g\left(\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)^{p^{\prime}}\mathrm{d}x\right)^{\frac{1}{p^{\prime}}},$
where now $p=\frac{L-s}{L}=1-\frac{s}{L}>0$ and $p^{\prime}=-(L-s)<0$. Since $g(x)^{p^{\prime}}\geq g(p^{\prime}x)$ for all $x$, we can lower-bound (21) further by
(22) $I_{M}(\lambda)^{1/p}\cdot\prod_{j=1}^{s}\left(\int_{0}^{1}\prod_{i=1}^{M}g\left(p^{\prime}\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)\mathrm{d}x\right)^{\frac{1}{p^{\prime}}}.$
For $|x|\leq 1$, we have $\exp(x)\geq 1+x+\frac{x^{2}}{2}+\frac{x^{3}}{6}>0$. By (10) we can indeed use this approximation to estimate $I_{M}(\lambda)$, and using an argument very similar to the one above (only without the quartic term, which is not needed now) we can show that
$I_{M}(\lambda)\geq\prod_{i=1}^{M}\left(1+\frac{\lambda^{2}p^{2}L}{2}+O(\lambda^{3}L)\right)$
as an analogue of (16). Since we are now interested in a lower bound, instead of $1+x\leq e^{x}$ we now use $1+x\geq e^{x-x^{2}}$, which is valid for $0\leq x\leq 1$. Note that we are allowed to use this inequality, since
$0\leq\frac{p^{2}\lambda^{2}L}{2}\leq\frac{4\lambda^{2}\frac{1}{4|\lambda|}}{2}\leq\frac{|\lambda|}{2}\leq 1$
by (5) and since $\lambda\in[-1,1]$. Thus we have
$I_{M}(\lambda)\geq\prod_{i=1}^{M}\exp\left(\frac{\lambda^{2}p^{2}L}{2}-\left(\frac{\lambda^{2}p^{2}L}{2}\right)^{2}+O(\lambda^{3}L)\right)=\exp\left(\frac{\lambda^{2}p^{2}LM}{2}+O(\lambda^{3}LM)\right).$
Note that the exponents $1/p^{\prime}$ are negative, so that for the integrals in (21) we again need an upper bound, and thus we can obtain the inequality
$\int_{0}^{1}\prod_{i=1}^{M}g\left(p^{\prime}\lambda\sqrt{{2}}\cos(2\pi n_{k_{i,j}}x)\right)\mathrm{d}x\leq 1$
in the same way as we did for (19). Overall, using (21) and the fact that $N=M(L+s)$, which implies $ML=N+O(\lambda N)$, we have
$\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x\geq$ $\exp\left(\frac{\lambda^{2}pLM}{2}+O(\lambda^{3}N)\right)$
$\geq$ $\exp\left(\frac{\lambda^{2}pN}{2}+O(\lambda^{3}N)\right).$
Together with (20), we have obtained matching upper and lower bounds, so that we finally arrive at
$\int_{0}^{1}\exp\left(\lambda\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)\right)\mathrm{d}x=\exp\left(\frac{\lambda^{2}pN}{2}+O(\lambda^{3}N)\right),$
as desired. ∎
Proof of Theorem 2.
We apply [13, Theorem 2.3.6] together with the subsequent remark. Let $t\in\mathbb{R}$ and $\lambda=\lambda_{N}$ be a null sequence with $\sqrt{N}\lambda\to\infty$. Then, by Theorem 1 it follows that
$\lim_{N\rightarrow\infty}\lambda^{2}\log\mathbb{E}\left[e^{t\frac{\lambda}{\sqrt{N}\lambda^{2}}\sum_{k=1}^{N}\sqrt{2}\cos(2\pi n_{k}x)}\right]$ $=\frac{t^{2}}{2}.$
The latter function is clearly essentially smooth and thus [13, Theorem 2.3.6] implies the claim by recalling that $t\mapsto\frac{t^{2}}{2}$ is a fixed point of the Legendre transform. ∎
Acknowledgments
CA is supported by the Austrian Science Fund (FWF), projects 10.55776/I4945, 10.55776/I5554, 10.55776/P34763 and 10.55776/P35322. MM is supported by the Austrian Science Fund (FWF), project 10.55776/DOC183.